# CancerGUIDE: Cancer Guideline Understanding via Internal Disagreement Estimation

### v1

@@ 4.1 Expert-Annotated Dataset
A total of 13 oncologists annotated 121 patient notes. To assess inter-annotator reliability, 11 examples were dually annotated, showing an average of 0.636 treatment match and 0.692 path overlap score. Notes were on average 54755.3 characters long and contained 82 distinct paths through the NCCN decision tree and 48 distinct final treatment recommendations. We use this annotation as ground truth for both evaluating model performance and validating proxy benchmarking approaches. By establishing a high-quality reference standard, we can interpret LLM performance, detect systematic errors, and calibrate models in high-stakes medical settings without relying exclusively on expensive ongoing human annotation.
@@
Table 1 illustrates the results of eight frontier models on the expert-annotated dataset. We report performance on GPT-5 (OpenAI, 2025b), GPT-4.1 (OpenAI, 2025a), o3 (OpenAI, 2025c), o4-mini (OpenAI, 2025d), DeepSeek-R1 (DeepSeek-AI, 2025), and LLaMA-3.3-70B-Instruct (Meta, 2024). We evaluate GPT-5 with varying reasoning efforts (minimal, medium, and high) to directly assess the impact of reasoning on performance. All models are evaluated with default temperature 1.0. We measured two clinically relevant metrics: (i) path overlap, measuring structural agreement with the annotated guideline path, and (ii) treatment match, indicating whether the recommended treatment node matches expert annotation.
Table 1: Model performance on Expert-Annotated Data
@@ 4.2 Proxy Benchmarks Enable Model Evaluation in Zero-Label Settings
We evaluated six proxy benchmarks for assessing LLM performance without human labels (Figure 2). Among synthetic approaches, introducing structure into the generation pipeline substantially increased error—RMSE rose from 0.11 (Synthetic Unstructured) to 0.43 (Synthetic Structured)—while providing only a negligible correlation gain (0.86 → 0.90). For consistency-based methods, thresholding by treatment match outperformed path overlap, likely because treatment match captures a more generalizable prediction goal, whereas path overlap harshly penalizes minor deviations with little impact on downstream treatment prediction. Aggregating outputs across models in cross-model consistency did not improve correlations and instead increased RMSE. Overall, synthetic benchmarks appear promising due to their scalability and robustness to variance between model consistency levels, but in domains where accuracy aligns with model consistency, consistency-based benchmarks can also provide a strong signal of performance. Examining model self-consistency in more detail reveals that outputs with higher internal agreement tend to be more accurate. We visualize this trend of iteration consistency in Figure 3 and report per-model Pearson correlation coefficients for both path overlap and treatment match in the appendix (Table 3), confirming that consistency is a strong indicator of instance-level accuracy.
@@ 4.3 Predicting Treatment Accuracy Using Model Self-Consistency
We develop a meta-classifier to predict when treatment recommendations are correct, using features from model consistency patterns and benchmark performance. Our approach achieves 0.800 AUROC by leveraging inference-time signals to predict accuracy without requiring ground truth labels at test time. We evaluate five feature sets to identify which signals contribute most to performance (Table 2). We stratify train and test sets by models and patient IDs with a 70/30 split. Performance generalizes across models and patients, enabling downstream classification with minimal human supervision.
@@
We find in Figure LABEL:fig:roc_aggregated that Base_aggregated, Aggregated_only, and All outperform other feature sets, indicating that cross-model consistency provides the strongest signal for accuracy classification. We see minimal gain from proxy benchmark performance, indicating consistency-based approaches are better suited for real-time analysis of model outputs than synthetic data. Even in settings without cross-model information (Base and Internal), we see high AUROC scores, indicating model self-consistency can be a strong signal for classification in isolation. Figure LABEL:fig:roc_base_agg_models shows intra-model variability of the classifier trained using Base_aggregated features, with results using Internal features reported in Appendix G. The strong performance of LLaMA-3.3-70B-Instruct highlights a key issue with including lower-quality models: its performance on the NCCN task is lower than other models, arbitrarily inflating AUROC as the confidence prediction task of the meta-classifier becomes easier. Besides LLaMA-3.3-70B-Instruct, GPT-4.1 has the highest AUROC and Deepseek-R1 the lowest, indicating GPT-4.1’s consistency better correlates with accuracy than Deepseek-R1.
Furthermore, in a fully unsupervised setting, we can separate accurate from inaccurate predictions with an F1 score of 0.666, demonstrating that consistency-based signals contain information for error detection without any labeled data.
@@ 5 Discussion
As LLMs become increasingly capable across tasks, understanding their limitations and quantifying accuracy for specific applications is critical for deployment (Unell et al., 2025). This work makes three main contributions in the high-stakes medical domain. First, we formalize guideline-concordant treatment recommendation generation and introduce a dataset of 121 patient cases annotated by board-certified oncologists. Second, we compare synthetic supervision and consistency-based supervision for benchmarking models on tasks lacking human annotations, finding that consistency-based benchmarks best approximate human judgments (Spearman $r=0.88$, RMSE=0.08). Third, we develop a hybrid agent and meta-classifier trained using unsupervised features, achieving an average of 0.800 AUC on held-out models and producing confidence scores that support ROC analysis, error trade-off optimization, and regulatory compliance (U.S. Food and Drug Administration, 2022). Feature analysis shows self-consistency and cross-model consistency are the strongest predictors of recommendation accuracy.
@@
Several promising directions emerge from this work. Expanding the dataset to other cancer types and guidelines would help characterize the generalizability of our consistency-based benchmarking approach, potentially enabling cross-domain accuracy prediction (Sorin et al., 2024; Pool and Trujillo, 2024). Further evaluation of self- and cross-model consistency as a signal for accuracy classification across broader model sizes and architectures could deepen understanding of this method’s robustness. Increasing both the size of the human-annotated dataset and the number of dually annotated samples would strengthen confidence in ground-truth labels. Since clinicians exhibit variability in path selection, future work should also explore ways to explicitly model and incorporate human uncertainty into evaluation. Adaptive learning approaches that account for uncertainty across clinical scenarios may further improve sensitivity of proxy metrics (Kompa et al., 2021; Jing et al., 2024). Finally, investigating how proxy benchmark data can be leveraged for alignment with downstream human preferences could help mitigate data bottlenecks that limit current methods (Kim and Seo, 2024; Lee et al., 2025).
@@ Appendix A Annotation Details
We begin by extracting the decision tree pages from the NCCN NSCLC guidelines (National Comprehensive Cancer Network, 2025) and creating a JSON-based representation where each decision point is labeled as a ”node.” Terminal nodes with ”recommendation” labels contain final treatment recommendations for clinicians, while intermediate nodes represent decision points requiring additional clinical information.
To evaluate model capability in mapping these guidelines onto real clinical data, we created a human-annotated dataset for guideline-adherent reasoning in NSCLC. Expert annotators traced patient notes through the NCCN guideline decision tree, recording the ordered sequence of page–node identifiers (e.g., NSCL-1-1 → NSCL-2-1 → … → terminal). Each annotation aimed to reach either the appropriate ”recommendation” node when sufficient clinical information was available, or the furthest node that could be accurately determined given the available patient data. Annotators additionally marked whether the observed care was guideline-compliant and provided brief rationales for any identified as non-compliance.
@@
A subset of cases was double-annotated, and disagreements were resolved through adjudication meetings, with all decisions documented for consistency. A strict no-AI-use policy was enforced via Insightful time-tracking with periodic screenshot audits (Insightful, 2025). Annotation time averaged 46.5 minutes per case. All cases were fully de-identified and contained no protected health information, so the study did not constitute human-subjects research and no IRB was required.
@@ Appendix C Treatment Match
$k$ = $=\text{ total rollouts}$ (1)
$c(t)$ $=\sum_{i=1}^{k}\mathbf{1}_{f_{i}=t}\quad\text{(count of paths with final treatment }t\text{)}$ (2)
The numerator $\max_{t\in T}c(t)$ counts the total number of ”repeated” final treatment occurrences (i.e., how many times each treatment appears beyond its first occurrence) and returns the frequency of the mode treatment selection. For all models $k=10$.
@@ Appendix D Error Analysis
We find that 40.42% of all errors made by models on human-annotated clinical text are flagged within the top 5 points of confusion by the model itself during its internal rollouts. This enables practitioners to refine instructions, perform fine-tuning, or apply other interventions to address these frequent errors—without relying on human annotations to detect them. Discrepancies mostly arise around guideline compliance and tumor staging, indicating that models struggle to differentiate between non-compliant and compliant cases with high accuracy as well as are unable to leverage parametric knowledge to perform tumor staging. Potential interventions could be training a model specifically for either task or providing additional context with clear instructions regarding clinical expectations.
@@ Appendix E Consistency-Accuracy Correlation
We show that for both Path Overlap metric and Treatment Match metrics, consistency of a prediction is correlated with accuracy. We do see, however, that there are different levels of correlation, indicating that consistency-based approaches inherently are more accurate for some model families than others. We note that DeepSeek-R1 has a low outlier correlation coefficient value. This highlights the robustness of the meta-classification pipeline to varying model internal consistencies. Even though there is no direct correlation between self-consistency and accuracy for this model, we are still able to classify accuracy with high fidelity in our method.
Table 3: Iteration consistency vs. accuracy correlation ($r$) for path overlap and treatment match. Mean $\pm$ SEM across models are reported at the bottom.
@@ Appendix F Unsupervised clustering performance
We cluster on features related to self- and cross-model consistency. We are able to distinguish between True Negatives and False Negatives with high accuracy, and are able to derive some signal regarding the True Positives. This indicates that consistency alone can be used as a baseline in an unsupervised method to identify accuracy of prediction. The performance is below supervised approaches, as to be expected, but is promising for further applications in which supervision is not viable or possible. F1 score is .666, indicating potential of unsupervised methods in conjunction with consistency approaches to evaluate model performance in zero-label settings.
@@ Appendix G Meta-Classifier Performance on Internal Features
We report the performance of our meta-classifier using internal features only for training. This enables us to disaggregate classification performance between cross-model and self consistency. We see here that the AUROC scores for each model is lower than when we include cross-model information, but it is still significantly above random, indicating that while cross-model consistency is a useful feature, it is not the only feature that allows for insight into model performance.
@@ Appendix H Model Information
For supervised classification, we trained a logistic regression classifier with L2 regularization (C=0.1), optimized using LBFGS with up to 10,000 iterations, and applied class-balanced weighting to account for label imbalance.

### v2

@@ 4.1 Expert-Annotated Dataset
A total of 13 oncologists annotated 121 patient notes, averaging 46.5 minutes per patient note with a cost of \$500 per US Board certified clinician hour. To assess inter-annotator reliability, 11 examples were dually annotated, showing an average of 0.636 treatment match and 0.692 path overlap score, with further disagreement analysis present in Appendix I. Notes were on average 54755 characters long and contained 82 distinct paths through the NCCN decision tree and 48 distinct final treatment recommendations. We use this annotation as ground truth for both evaluating model performance and validating proxy benchmarking approaches. By establishing a high-quality reference standard, we can interpret LLM performance, detect systematic errors, and calibrate models in high-stakes medical settings without relying exclusively on expensive ongoing human annotation.
@@
Table 1 illustrates the results of eight frontier models on the expert-annotated dataset. We report performance on GPT-5 (openai2025gpt5), GPT-4.1 (openai2025gpt41), o3 (openai2025o3), o4-mini (openai2025o4mini), DeepSeek-R1 (deepseek2025r1), and LLaMA-3.3-70B-Instruct (meta2024llama3). We evaluate GPT-5 with varying reasoning efforts (minimal, medium, and high) to directly assess the impact of reasoning on performance. All models are evaluated with default temperature 1.0. We measured two clinically relevant metrics: (i) path overlap, measuring structural agreement with the annotated guideline path, and (ii) treatment match, indicating whether the recommended treatment node matches expert annotation.
Table 1: Model performance on Expert-Annotated Data (mean $\pm$ SEM).
@@ 4.2 Proxy Benchmarks Enable Model Evaluation in Zero-Label Settings
We evaluated six proxy benchmarks for assessing LLM performance without human labels (Figure 2). Among synthetic approaches, adding structure to the generation pipeline substantially increased error (RMSE rising from 0.11 in Synthetic Unstructured to 0.43 in Synthetic Structured) while yielding only a marginal correlation gain (Spearman $r$: 0.86 → 0.90). For consistency-based methods, thresholding by treatment match outperformed path overlap, likely because treatment match reflects a more generalizable prediction goal, whereas path overlap penalizes minor deviations that have little effect on downstream treatment accuracy. Aggregating outputs across models in cross-model consistency failed to improve correlations and instead increased RMSE. Overall, synthetic benchmarks are appealing for their robustness to variation in model consistency, while consistency-based benchmarks are strong in domains where accuracy and consistency are correlated. In this setting, self-consistency is aligned with model accuracy (Figure 3), and per-model Pearson correlations reported in Appendix E further confirm that consistency reliably indicates instance-level accuracy.
@@ 4.3 Predicting Treatment Accuracy Using Model Self-Consistency
We develop a meta-classifier to predict when treatment recommendations are correct, using features from model consistency patterns and benchmark performance. Our approach achieves 0.804 AUROC by leveraging inference-time signals to predict accuracy without requiring ground truth labels at test time. We evaluate five feature sets to identify which signals contribute most to performance (Table 2). We stratify train and test sets of the human annotated benchmark by models and patient IDs with a 70/30 split. Performance generalizes across models and patients, enabling downstream classification with minimal human supervision.
@@
We find in Figure LABEL:fig:roc_aggregated that Base_aggregated, Aggregated_only, and All outperform other feature sets, indicating that cross-model consistency provides the strongest signal for accuracy classification. We see minimal gain from proxy benchmark performance, indicating consistency-based approaches are better suited for real-time analysis of model outputs than synthetic data. Even in settings without cross-model information (Base and Internal), we see high AUROC scores, indicating model self-consistency can be a strong signal for classification in isolation. Figure LABEL:fig:roc_base_agg_models shows intra-model variability of the classifier trained using Base_aggregated features, with results using Internal features reported in Appendix G to highlight that even in cases where consistency and accuracy are not strongly correlated, such as for DeepSeek-R1, these features still provide some signal for accuracy classification with an AUROC of 0.582. The strong classification performance of LLaMA-3.3-70B-Instruct outputs highlights a key issue with including lower-quality models: its performance on the NCCN task is lower than other models, arbitrarily inflating AUROC as the confidence prediction task of the meta-classifier becomes easier. Besides LLaMA-3.3-70B-Instruct, GPT-4.1 has the highest AUROC and Deepseek-R1 the lowest, indicating GPT-4.1’s consistency better correlates with accuracy than Deepseek-R1.
Furthermore, in a fully unsupervised setting, we are able to separate accurate from inaccurate predictions with an F1 score of 0.666 compared to the average F1 score of 0.702 in a supervised setting, demonstrating that consistency-based signals carry meaningful information for error detection without any labeled data (F1 scores for supervised classification are thresholded with value 0.5). Notably, 40.42% of all model errors on human-labeled data can be detected without relying on any human labels (detailed analysis in Appendix D). These results indicate that consistency signals alone provide sufficient structure to distinguish between accurate and inaccurate predictions and to identify specific failure modes, enabling model developers to detect and address potential issues prior to deployment and support iterative refinement of clinical decision support systems.
@@ 5 Discussion
As LLMs become increasingly capable across tasks, understanding their limitations and quantifying accuracy for specific applications is critical for deployment (realworld). This work makes three main contributions in the high-stakes medical domain. First, we formalize guideline-concordant treatment recommendation generation and introduce a dataset of 121 patient cases annotated by board-certified oncologists. Second, we compare synthetic supervision and consistency-based supervision for benchmarking models on tasks lacking human annotations, finding that consistency-based benchmarks best approximate human judgments (Spearman $r=0.88$, RMSE=0.08). Third, we develop a hybrid agent and meta-classifier trained using unsupervised features, achieving an average of 0.804 AUC on held-out models and producing confidence scores that support ROC analysis, error trade-off optimization, and regulatory compliance (fda). ROC curves allow end-users—clinicians or developers—to adjust thresholds to balance sensitivity and specificity according to context, aligning with FDA guidance recommending reporting of ROC curves and related metrics, since the relative costs of false positives and false negatives vary by diagnostic scenario. Importantly, calibrated confidence scores derived from our meta-classifier provide clinicians with interpretable indicators of reliability, enabling informed decisions about the weight to assign to generated recommendations in high-stakes settings. For example, imagine an oncologist needing to recommend a second-line therapy for a patient with NSCLC. A calibrated confidence score allows the clinician to assess the reliability of the model’s guideline-adherent recommendation, which can serve as a baseline option. The clinician can then decide whether to proceed with this recommendation or intentionally deviate toward a personalized, non-guideline based approach—now with an informed understanding of the default, evidence-based standard of care. Our analysis also demonstrates that consistency signals can reveal model failure modes—such as difficulty detecting guideline non-compliance or TNM staging errors—even without extensive human annotation, highlighting both opportunities for targeted model improvement and the practical utility of these metrics in clinical contexts. Feature analysis shows that these consistency signals (self-consistency and cross-model consistency) are additionally the strongest predictors of model recommendation accuracy. This supports that consistency signals can be utilized both for identifying errors in a label-free setting and also for calibrating confidence to allow users to make informed decisions regarding their trust in the system output.
@@
Several promising directions emerge from this work. Expanding the dataset to other cancer types and guidelines would help characterize the generalizability of our consistency-based benchmarking approach, potentially enabling cross-domain accuracy prediction (cancer1; cancer2). Further evaluation of self- and cross-model consistency as a signal for accuracy classification across broader model sizes and architectures could deepen understanding of this method’s robustness. Increasing both the size of the human-annotated dataset and the number of dually annotated samples would strengthen confidence in ground-truth labels. Since clinicians exhibit variability in path selection, future work should also explore ways to explicitly model and incorporate human uncertainty into evaluation. Adaptive learning approaches that account for uncertainty across clinical scenarios may further improve sensitivity of proxy metrics (active1; active2). Finally, investigating how proxy benchmark data can be leveraged for alignment with downstream human preferences could help mitigate data bottlenecks that limit current methods (proxy1; proxy2).
@@ Appendix A Annotation Details
We begin by extracting the decision tree pages from the NCCN NSCLC guidelines (NCCN) and creating a JSON-based representation where each decision point is labeled as a “node.” Terminal nodes with “recommendation” labels contain final treatment recommendations for clinicians, while intermediate nodes represent decision points requiring additional clinical information.
To evaluate model capability in mapping these guidelines onto real clinical data, we created a human-annotated dataset for guideline-adherent reasoning in NSCLC. Expert annotators traced patient notes through the NCCN guideline decision tree, recording the ordered sequence of page–node identifiers (e.g., NSCL-1-1 → NSCL-2-1 → … → terminal). Each annotation aimed to reach either the appropriate “recommendation” node when sufficient clinical information was available, or the furthest node that could be accurately determined given the available patient data. Annotators additionally marked whether the observed care was guideline-compliant and provided brief rationales for any identified as non-compliance. Non-compliance is treated as a valid “path” and evaluates model’s capacity to identify patient treatment trajectories as non-compliant with given guidelines.
@@
A subset of cases was double-annotated, and disagreements were resolved through adjudication meetings, with all decisions documented for consistency. A strict no-AI-use policy was enforced via Insightful time-tracking with periodic screenshot audits (insightful). Annotation time averaged 46.5 minutes per case. All cases were fully de-identified and contained no protected health information, so the study did not constitute human-subjects research and no IRB was required.
@@ Appendix C Treatment Match
$k=\text{total rollouts}$
$c(t)=\sum_{i=1}^{k}\mathbf{1}_{\{f_{i}=t\}}$
The numerator $\max_{t\in T}c(t)$ counts the total number of “repeated” final treatment occurrences (i.e., how many times each treatment appears beyond its first occurrence) and returns the frequency of the mode treatment selection. For all models $k=10$.
@@ Appendix D Error Analysis
We find that 40.42% of all errors made by models on human-annotated clinical text are flagged within the top 5 points of confusion by the model itself during its internal rollouts. This enables practitioners to refine instructions, perform fine-tuning, or apply other interventions to address these frequent errors—without relying on human annotations to detect them. Figure LABEL:fig:ea_comparison shows the overlap between errors identified by model consistency vs. the true errors made by the model against human annotated data. Discrepancies mostly arise around guideline compliance and tumor staging, indicating that models struggle to differentiate between non-compliant and compliant cases with high accuracy as well as are unable to leverage parametric knowledge to perform tumor staging. Potential interventions could be training a model specifically for either task or providing additional context with clear instructions regarding clinical expectations.
@@ Appendix E Consistency-Accuracy Correlation
We show that for both Path Overlap metric and Treatment Match metrics, consistency of a prediction is correlated with accuracy (Table 3). We do see, however, that there are different levels of correlation, indicating that consistency-based approaches inherently are more accurate for some model families than others. We note that DeepSeek-R1 has an outlier correlation coefficient value, being negative for Path Overlap score and extremely low for Treatment Match score’s correlation with accuracy. This highlights the robustness of the meta-classification pipeline to varying model internal consistencies. Despite the lack of direct correlation between self-consistency and accuracy for this model, we are still able to classify accuracy with high fidelity.
Table 3: Iteration consistency vs. accuracy correlation ($r$) for path overlap score and treatment match score. Mean correlation values across all models $\pm$ SEM are reported at the bottom.
@@ Appendix F Unsupervised clustering performance
We cluster on features related to self- and cross-model consistency (Figure 15). We are able to distinguish between True Negatives and False Negatives with high accuracy, and are able to derive some signal regarding the True Positives. This indicates that consistency alone can be used as a baseline in an unsupervised method to identify accuracy of prediction. The performance is below supervised approaches, as to be expected, but is promising for further applications in which supervision is not viable or possible. F1 score is .666, indicating potential of unsupervised methods in conjunction with consistency approaches to evaluate model performance in zero-label settings.
@@ Appendix G Meta-Classifier Performance on Internal Features
We report the performance of our meta-classifier using internal features only for training, as shown in Figure 16. This enables us to disaggregate classification performance between cross-model and self consistency. We see here that the AUROC scores for each model is lower than when we include cross-model information, but it is still significantly above random, indicating that while cross-model consistency is a useful feature, it is not the only feature that allows for insight into model performance.
@@ Appendix H Model Information
For supervised classification, we trained a logistic regression classifier with L2 regularization (C=0.1), optimized using LBFGS with up to 10,000 iterations, and applied class-balanced weighting to account for label imbalance. Table 4 reports further API details.
@@
Appendix I Human Annotation Disagreement Analysis
Of the 11 dually annotated records:
  1. 1.
    7 were agreed upon by all physicians.
  2. 2.
    4 required adjudication
  3. 3.
    2 involved disagreements about whether the patient’s history complied with guidelines (these cases are excluded from our benchmark).
  4. 4.
    1 involved assumptions regarding the patient’s history; after clarification to rely strictly on the provided record, both clinicians agreed.
  5. 5.
    1 involved a TNM staging error by one clinician, resolved upon discussion.
After removing non-compliant cases, agreement across the benchmark was 7/9. This reflects the inherent difficulty and nuance of the task, setting a realistic upper bound for LLM performance in clinical decision support scenarios.
Appendix J NCCN Tree Generation
To construct the clinical decision tree, we first manually structured a subset of sample guideline pages into a JSON-based hierarchical format representing decision nodes and treatment branches. These manually curated examples served as few-shot demonstrations for prompting the language model to generate additional decision trees for the remaining guideline sections. The generated JSON outputs were automatically validated to ensure structural integrity and adherence to the required schema. While this process provided an initial level of validation, the resulting trees were not independently verified by clinical experts, and thus may contain minor inconsistencies or omissions relative to expert-curated guideline representations. The generated NCCN tree for NSCLC has 3,302 nodes, 3,301 edges, and a max depth of 8.
Appendix K Clinical Data Details
The patient data used for analysis comprised approximately 2,900 patients diagnosed with breast and lung cancer at various stages, drawn from two U.S. health systems: a community-based system and a large nonprofit health system, each spanning multiple hospital locations. The dataset included electronic medical records, radiology reports, pathology reports, and other clinical documents such as next-generation sequencing (NGS) reports. Tumor registry data were obtained from three complementary sources: (1) the hospitals’ internal registries, (2) registry submissions to the state (for one system), and (3) manually labeled registry data extracted by over 30 registered oncology nurses. All data were de-identified by the vendor prior to delivery, following HIPAA-compliant procedures that included redaction of all PHI, date shifting relative to the date of birth, and removal of DICOM headers and pathology metadata containing identifiers.